A body is just floating in a liquid ( their densities are equal ). What will happen to the body if it is slightly pressed and released ? Why ?

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Assuming Archimedes' principle to be reformulated as follows,

\mbox{apparent immersed weight} = \mbox{weight} - \mbox{weight of displaced fluid}\,

then inserted into the quotient of weights, which has been expanded by the mutual volume

 \frac { \mbox{density}} { \mbox {density of fluid} } = \frac { \mbox{weight}} { \mbox {weight of displaced fluid} } \,,

yields the formula below. The density of the immersed object relative to the density of the fluid can easily be calculated without measuring any volumes:

 \frac { \mbox {density of object}} { \mbox {density of fluid} } = \frac { \mbox {weight}} { \mbox {weight} - \mbox {apparent immersed weight}}\,

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi